OpenStudy (anonymous):

i need help finding the percentage change from 82 to 38

hero (hero):

\[\% \Delta = \frac{|a-b|}{a} \times 100\] a = original number b = new number

hero (hero):

\[\%\Delta = \frac{82 - 38}{82} \times 100\]

hero (hero):

\[\% \Delta = 53.6585\]

OpenStudy (anonymous):

so which one is the original number

hero (hero):

The first number you read is the original number

hero (hero):

Think of the original question as asking: What is the percent change from a to b?

hero (hero):

This would mean that 82 = a , 38 = b

OpenStudy (anonymous):

so you dont cross multiply

hero (hero):

All you do is input the values for a and b into the formula and solve.

OpenStudy (anonymous):

ok so i get confused dont knw what number to divide

OpenStudy (anonymous):

thanks anyway